# Initial-state geometry and multiplicity distributions in pp and pPb collisions

### v2

@@ Abstract
This work investigates the possibility of accessing the initial geometric shape of the proton in proton-proton and proton-nucleus collisions at the LHC, in particular the configuration in which the proton is made of three quarks linked by a Y-shape gluon string, called baryon junction. This initial state spatial configuration has been used in the past to describe data on baryon rapidity distributions, diffractive $J/\psi$ production and multiplicity distributions in pp collisions. In spite of its success in explaining the data, the evidence of the baryon junction still needs confirmation. Further studies will be undertaken at the electron-ion collider. In this work we study multiplicity distributions measured in pp and pPb collisions. Different initial state geometries are used as input in a Monte Carlo event generator which implements the $k_{T}$-factorization formalism of the CGC with KLN unintegrated gluon distributions. The results indicate that the hard-sphere and Gaussian proton configurations are incompatible with the data. In contrast, the baryon junction configuration can describe the data provided that intrinsic fluctuations of the saturation scale are included.
@@ I.1 The baryon junction
Fortunately, there are good reasons to believe that the Y-shape configuration is not washed out and manifests itself even in high energy collisions. Indeed, there are indications that the geometric organization of matter persists even at LHC energies. In manty it was shown that exclusive vector meson production are sensitive to the geometric deformation of the target. The authors concluded that the nuclear geometry (and its deformations) is not washed out by small-x evolution and hence the deformations previously inferred from low-energy experiments will be visible in high-energy collisions. This suggests the Y shape of the proton could survive the quantum evolution and manifest itself in high energy pp collisions. In view of the uncertainties involving this issue, more calculations and more data are needed. Along this direction, there is an interesting proposal to look for the BJ at the forthcoming electron-ion collider bran22 .
@@ I.2 The baryon junction and multiplicity distributions
From the point of view of modelling, the initial state geometry can be naturally implemented in the partonic version of the Glauber model glauber07 ; loi16 ; loi19 . Typically one chooses the initial spatial density of matter and, with it, computes the thickness function, which appears in the expressions of the cross sections. In this work we will simulate the collisions in the MC-KLN event generator mckln . This Monte Carlo simulator implements the $k_{T}$-factorization formalism of the Color Glass Condensate (CGC) using the KLN unintegrated gluon distribution kln1 ; kln2 . For a nucleon with a given initial spatial configuration, this simulator allows one to compute the gluon production. Conversions to final-state hadrons are taken into account via the parton-hadron duality, meaning that the particle production at parton and hadron level only differs by a constant factor, also employed in previous phenomenlogical studies involving $p_{T}$-integrated observables calculated via the $k_{T}$-factorization formalism mckln ; kln1 ; kln2 ; Dumitru:2011wq ; Dumitru:2012yr ; Albacete:2012xq ; Duraes:2016yyg . With these tools we can use the baryon junction initial conditions and calculate the final multiplicity distribution, looking for its manifestations.
@@ II.1 Initial conditions for the nucleon
We start by introducing four types of nucleon matter spatial distributions used in our simulations: hard-sphere, Gaussian, analytical baryon junction (from now on called BJ1) and a slightly different and purely numerical baryon junction (from now on called BJ2). Examples of the thickness function (z-integrated density) of each configuration are shown on the left panel of Fig. 1. The configurations are described below.
@@
Figure 1: (Left) Proton thickness. (Middle) Lead participant thickness. (Right) $dN/dy\,d^{2}\mathbf{r}$ of the produced gluons at midrapidity. The rows represent different density distributions of the nucleons. From top to bottom: hard-sphere, Gaussian, BJ1, BJ2.
@@ II.1.1 Hard-sphere
The hard-sphere is based on the collision criterion of distances between two nucleon centers. Explicitly, there is a collision when the distance between the projectile (proj) and the target (targ) is:
$d^{2}=\big(x_{proj}-x_{targ}\big)^{2}+\big(y_{proj}-y_{targ}\big)^{2}\leq\frac{\sigma_{NN}}{\pi}\,$, (1)
where $\sigma_{NN}$ is the nucleon-nucleon inelastic cross section for a given center of mass energy. This collision criterion implies that each nucleon is geometrically described by a filled disk, with thickness given by:
[…]
@@ II.1.2 Gaussian nucleon
with $B_{p}$ being the nucleon width, which is related to the root mean square radius of the proton as $r_{p}=\sqrt{2B_{p}}$. Older fits to HERA data reza , as well as recent Bayesian analysis based solely on information from $e+p$ collisions from HERA Mantysaari:2022ffw , show that $B_{p}\approx 4\,{\rm GeV}^{2}$, while recent Bayesian analysis of combined data on diffractive $J/\Psi$ photoproduction in $\gamma+p$ and $\gamma+Pb$ collisions yield smaller values Mantysaari:2025ltq .
@@
$P(b)=1-\exp\big(-\sigma^{eff}T_{NN^{\prime}}(b)\big)\,,\qquad T_{NN^{\prime}}(b)=\int T_{N}(x-b/2,y)T_{N^{\prime}}(x+b/2,y)dxdy=\frac{e^{-b^{2}/4B}}{4\pi B}\,$, (4)
@@ II.1.3 Baryon junction BJ1
$T_{N}^{q}(x,y)=\frac{(1-\kappa)}{6\pi B_{p}^{2}}\sum_{i=1}^{3}\exp\Bigg(-\frac{\big(x-(x_{p}+x_{i})\big)^{2}+\big(y-(y_{p}+y_{i})\big)^{2}}{2B_{p}^{2}}\Bigg)\,$, (6)
@@
$T_{N}^{g}(x,y)=\frac{\kappa}{6\pi B_{p}^{2}}\sum_{i=1}^{3}\exp\Bigg(-\frac{\big(x-(x_{p}+\frac{x_{i}}{2})\big)^{2}+\big(y-(y_{p}+\frac{y_{i}}{2})\big)^{2}}{2B_{p}^{2}}\Bigg)\,$, (7)
In the formulas above, $\mathbf{r}_{p}=(x_{p},y_{p})$ is the coordinate of the center of the nucleon in the transverse plane, $\kappa=0.5$ is the gluon fraction, $B_{p}=0.17\text{ fm}$ is the hot-spot size. All these parameters were fixed in previous works nosso ; deb20 ; schenke ; disserta ; glazek-p .
@@
$=$ $\frac{(1-\kappa)^{2}}{36\pi B_{p}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i})-(x_{a}+x_{j})\big)^{2}+\big((y_{p}+y_{i})-(y_{a}+y_{j})\big)^{2}}{4B_{p}^{2}}\bigg)\,$,
@@
$=$ $\frac{\kappa^{2}}{36\pi B_{p}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\Bigg(-\frac{\big(b+(x_{p}+x_{i}/2)-(x_{a}+x_{j}/2)\big)^{2}+\big((y_{p}+y_{i}/2)-(y_{a}+y_{j}/2)\big)^{2}}{4B_{p}^{2}}\Bigg)\,$,
@@
$=$ $\frac{\kappa(1-\kappa)}{36\pi B_{p}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i}/2)-(x_{a}+x_{j})\big)^{2}+\big((y_{p}+y_{i}/2)-(y_{a}+y_{j})\big)^{2}}{8B_{p}^{2}}\bigg)\,$,
@@
$=$ $\frac{\kappa(1-\kappa)}{36\pi B_{p}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i})-(x_{a}+x_{j}/2)\big)^{2}+\big((y_{p}+y_{i})-(y_{a}+y_{j}/2)\big)^{2}}{8B_{p}^{2}}\bigg)\,.$
@@ II.1.4 Baryon Junction BJ2
The baryon junction BJ2 is a numerical Ansatz for the Y-like structure, as imployed in schenke . In this approach we draw the position of the three quarks in the 3D space according to the distribution:
@@ II.2 Initial condition for the nucleus
where $\rho_{0}$ is the nucleon density in the center of the nucleus, $R$ is the radius of the nucleus, $w=0$ represents deviations from the sphere format, and $a$ is the skin depth. Some usual values for these parameters can be found in Pbparam . We perform our simulation by sorting 208 nucleon center positions, $\mathbf{r}_{n}$, and then inserting the chosen nucleon distribution around it. In the end, the Pb thickness is:
@@ II.3 $k_{T}$-factorization
$\phi_{KLN}(\mathbf{k},x)=\frac{2C_{F}}{3\pi^{2}}\frac{(1-x)^{4}}{\alpha_{s}}\left\{\begin{matrix}1,&\text{, if }k\leq Q_{s}\\ \big(Q_{s}/k\big)^{2}&\text{, if }k>Q_{s}\end{matrix}\right.$. (19)
@@
$Q_{s}^{2}(x,\mathbf{r_{\perp}})=T(\mathbf{r_{\perp}})\,\,\frac{2\text{ GeV}^{-2}}{1.53\text{ fm}^{-2}}\,\,\bigg(\frac{0.01}{x}\bigg)^{\bar{\lambda}}\,$, (20)
@@
The MC-KLN convolutes the initial conditions of p and p (Pb) with the $k_{T}$-factorization formula of Eq.(17) by calculating $dN/dy\,d^{2}\mathbf{r_{\perp}}$ of gluons at a given rapidity in the overlap region. One example of each initial state geometry for pPb collisions is shown in Fig. 1, for the same seed. On the left column we present the proton thickness, on the middle column we show the thickness of the lead nucleus and on the right column we show the convolution of the initial conditions and Eq.(17) at midrapidity. We can observe that, remarkably, the produced gluon distribution on the right column inherits the proton shape, which makes pPb collisions good systems to probe the proton shape. It happens because the proton size, which is about $1\text{ fm}$, is much smaller than the lead radius ($6.62\text{ fm}$). Thus, the overlap region will invariably present the proton shape. On the other hand, in pp collisions, since the projectile and target are comparable in size, only fragments of the proton structure will overlap, which makes this system less sensitive to the entire structure.
@@ II.4 Intrinsic fluctuations
So far we have considered just geometrical fluctuations, i.e., those arising from changing positions of the initial quarks and gluons in the transverse plane. In order to describe high-multiplicity events, we must include dynamical fluctuations in the saturation scale larry16 of a given nucleon in the collision. These fluctuations are characterized by deviations of the saturation scale from its mean value and they follow the distribution:
@@ III Results
Integrating Eq.(17) in $p_{T}$, in $y$ (or $\eta$) and in $r_{\perp}$ in an event-by-event basis, allows one to construct the probablity $P(N_{ch})$ of having $N_{ch}$ charged particles produced. The results for pp and pPb collisions are shown in Fig. 2 and Fig. 3, respectively. In the figures $K$ is the multiplicative constant of Eq.(17), adjusted in order to reproduce the measured $\left\langle N_{ch}\right\rangle$. The parameter $\sigma$ is the dynamical fluctuation constant in Eq.(22). In Figs. 3b and 3d we see the crucial role played by intrinsic fluctuations at large values of $N_{ch}$.
Results for $P(N_{ch})$ are often presented in terms of the so called Koba-Nielsen-Olsen (KNO) scaling Koba:1972ng , which states that the multiplicity of produced particles becomes energy independent at asymptotically high energies when scaling the $N_{ch}$ by its mean value, $\langle N_{ch}\rangle$. This exact feature has been verified experimentaly at the LHC for narrow pseudo-rapidity ($|\eta|<0.5$) intervals CMS:2010qvf ; at the same time, the violation of this scaling has been reported for larger pseudo-rapidity intervals.
In order to reconcile the CGC formalism with KNO scaling, it has been shown that the theory must be nearly Gaussian (corresponding to the strong color field limit) and running coupling effects must be included as well Dumitru:2012tw . Such result was only verified for pp collisions up to $\sim 3.5$ times the average multiplicity. Small violations of the KNO scaling have been reported for p+Pb collisions in the region where $N_{ch}\lesssim 3$, where additional sources of fluctuations (geometric) are present Dumitru:2012yr . Here, we extend the investigation to larger multiplicities. It is worth noting that KNO scaling might need to be modified for very large multiplicity events Moriggi:2025qfs .
In spite of all these efforts, a clear connection between KNO scaling and QCD is not yet established. For complemetary references on this subject we refer the reader to henrique ; germano1 ; germano2 .
The first measurements of multiplicity distribution in pPb collisions alice24 were taken at two different energies. It is possible to make a first check of the validity of the KNO, which, as can be seen in Fig. 4a, seems to be valid. In that figure we see that both experimental data and the MC-KLN model with BJ2 initial geometry exhibit KNO scaling.
Looking at these figures we conclude that these data rule out the hard-sphere and the Gaussian models for pp, and also BJ1 for pPb. The BJ2 reproduces the data well for $N_{ch}>\left\langle N_{ch}\right\rangle$ in both configurations. Another notable characteristic is that the results from BJ1 and BJ2 are almost the same for pp, but very different for pPb. This is a consequence of the high sensitivity of pPb to the proton’s initial condition. In other words, pp collisions do not distinguish the numerical baryon junction and the analytical one. In the low multiplicity region almost all curves slightly underpredict the data. This may be due to the fact that in our formalsim we did not include the contribution from the valence quarks. They are supposed to be important only in the large rapidity regions, where one can use the hybrid formalism duja . Here, at lower rapidities their contribution should be less important but still needed for a better agreement with data.
@@
We can use the BJ2 initial condition to estimate the multiplicity distributions for different pseudorapidity bins. In Fig.4b we show our predictions for the multiplicity distribution in the pseudorapidity range $|\eta|<0.5$ and $|\eta|<1.0$. To do that we repeat the same calculation as before, with the same $K=1.84$ and $\sigma=0.4$ constant. We also provide the mean value estimate and observe that KNO scaling is still valid for these pseudorapidity intervals.
IV Conclusion
This work is part of a broader project of looking for experimental manifestations of the baryon junction configuration of the proton. Here we investigated proton-proton and proton-nucleus collisions at the LHC. More specifically, we calculated the multiplicity distributions measured in pp and pPb collisions. We used different initial state geometries as input in the MC-KLN Monte Carlo event generator which implements the $k_{T}$-factorization formalism of the CGC with KLN unintegrated gluon distributions. We found that the hard-sphere and the Gaussian proton configurations are incompatible with pp data, and the BJ1 is also incompatible with pPb. The baryon junction configuration BJ2 can describe the data provided that fluctuations in the saturation scale (also called “intrinsic fluctuations”) are included.
@@ IV Conclusion
Figure 4: a) KNO plot for $\sqrt{s}=5.02,8.16\text{ TeV}$. The experimental data are from alice24 ; b) Predictions for pPb multiplicity distributions at $\sqrt{s}=5.02\text{ TeV}$ for two rapidity intervals $|\eta|<0.5$ and $|\eta|<1.0$ presented in the KNO form.

### v3

@@ Abstract
This work investigates the possibility of accessing the initial geometric shape of the proton in proton-proton and proton-nucleus collisions at the LHC. In particular, we look for manifestations of the configuration in which the proton is made of three quarks linked by a Y-shape gluon string, called baryon junction. This initial state spatial configuration has been used in the past to describe data on baryon rapidity distributions, diffractive $J/\psi$ production and multiplicity distributions in pp collisions. In spite of its success in explaining the data, the evidence of the baryon junction still needs confirmation. Further studies will be undertaken at the electron-ion collider. In this work we study multiplicity distributions measured in pp and pPb collisions. Different initial state geometries are used as input in a Monte Carlo event generator which implements the $k_{T}$-factorization formalism of the CGC with KLN unintegrated gluon distributions. The results show that the data on multiplicity distributions are good enough to discriminate between different initial state geometries. Moreover, they indicate that it is crucial to take into account the intrinsic fluctuations of the saturation scale.
@@ I.1 The baryon junction
Fortunately, there are good reasons to believe that the Y-shape configuration is not washed out and manifests itself even in high energy collisions. Indeed, there are indications that the geometric organization of matter persists even at LHC energies. In manty it was shown that exclusive vector meson production is sensitive to the geometric deformation of the target. The authors concluded that the nuclear geometry (and its deformations) is not washed out by small-x evolution and hence the deformations previously inferred from low-energy experiments will be visible in high-energy collisions. This suggests the Y shape of the proton could survive the quantum evolution and manifest itself in high energy pp collisions. In view of the uncertainties involving this issue, more calculations and more data are needed. Along this direction, there is an interesting proposal to look for the BJ at the forthcoming electron-ion collider bran22 .
@@ I.2 The baryon junction and multiplicity distributions
From the point of view of modelling, the initial state geometry can be naturally implemented in the partonic version of the Glauber model glauber07 ; loi16 ; loi19 . Typically one chooses the initial spatial density of matter and, with it, computes the thickness function, which appears in the expressions of the cross sections. In this work we will simulate the collisions in the MC-KLN event generator mckln . This Monte Carlo simulator implements the $k_{T}$-factorization formalism of the Color Glass Condensate (CGC) using the KLN unintegrated gluon distribution kln1 ; kln2 . For a nucleon with a given initial spatial configuration, this simulator allows one to compute gluon production. Conversions to final-state hadrons are taken into account via the parton-hadron duality, meaning that the particle production at parton and hadron level only differ by a constant factor, also employed in previous phenomenlogical studies involving $p_{T}$-integrated observables calculated via the $k_{T}$-factorization formalism mckln ; kln1 ; kln2 ; Dumitru:2011wq ; Dumitru:2012yr ; Albacete:2012xq ; Duraes:2016yyg . With these tools we can use the baryon junction initial conditions and calculate the final multiplicity distribution, looking for its manifestations.
@@ II.1 Initial conditions for the nucleon
We start by introducing four types of nucleon matter spatial distributions used in our simulations: hard-sphere, Gaussian, analytical baryon junction (from now on called BJ1) and a slightly different and purely numerical baryon junction (from now on called BJ2). Examples of the thickness function (z-integrated density) of each configuration are shown on the left panels of Figs. 1 and 2. The configurations are described below.
@@
Figure 1: Proton-proton collisions. (Left) Projectile proton thickness. (Middle) Target proton thickness. (Right) $dN/dy\,d^{2}\mathbf{r}$ of the produced gluons at midrapidity. The rows represent different density distributions of the nucleons. From top to bottom: hard-sphere, Gaussian, BJ1, BJ2.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Proton-lead collisions. (Left) Proton thickness. (Middle) Lead participant thickness. (Right) $dN/dy\,d^{2}\mathbf{r}$ of the produced gluons at midrapidity. The rows represent different density distributions of the nucleons. From top to bottom: hard-sphere, Gaussian, BJ1, BJ2.
@@ II.1.1 Hard-sphere
The hard-sphere is based on the collision criterion of distances between two nucleon centers Albacete:2012xq . Explicitly, there is a collision when the distance between the projectile (proj) and the target (targ) is:
$d^{2}=\big(x_{proj}-x_{targ}\big)^{2}+\big(y_{proj}-y_{targ}\big)^{2}\leq\frac{\sigma_{0}}{\pi}\,$, (1)
where $\sigma_{0}$ is the transverse area of the large $x$ partons of a nucleon. As in Ref. Albacete:2012xq , we choose the parameter $\sigma_{0}$ to be equal to $42$ mb. This collision criterion implies that each nucleon is geometrically described by a filled disk, with thickness given by:
$T(\mathbf{r}_{\perp})=\frac{\Theta(\sqrt{\sigma_{0}/\pi}-r_{\perp})}{\sigma_{0}}\,$, (2)
@@ II.1.2 Gaussian nucleon
with $B_{p}$ being the nucleon width, which is related to the root mean square radius of the proton as $r_{p}=\sqrt{2B_{p}}$. Older fits to HERA data reza , as well as a recent Bayesian analysis based solely on information from $e+p$ collisions from HERA Mantysaari:2022ffw , show that $B_{p}\approx 4\,{\rm GeV}^{-2}$, while a recent Bayesian analysis of combined data on diffractive $J/\Psi$ photoproduction in $\gamma+p$ and $\gamma+Pb$ collisions yields a smaller value Mantysaari:2025ltq .
@@
$P(b)=1-\exp\big(-\sigma^{eff}T_{NN^{\prime}}(b)\big)\,,\qquad T_{NN^{\prime}}(b)=\int T_{N}(x-b/2,y)T_{N^{\prime}}(x+b/2,y)dxdy=\frac{e^{-b^{2}/4B_{p}}}{4\pi B_{p}}\,$, (4)
@@ II.1.3 Baryon junction BJ1
$T_{N}^{q}(x,y)=\frac{(1-\kappa)}{6\pi B_{h}^{2}}\sum_{i=1}^{3}\exp\Bigg(-\frac{\big(x-(x_{p}+x_{i})\big)^{2}+\big(y-(y_{p}+y_{i})\big)^{2}}{2B_{h}^{2}}\Bigg)\,$, (6)
@@
$T_{N}^{g}(x,y)=\frac{\kappa}{6\pi B_{h}^{2}}\sum_{i=1}^{3}\exp\Bigg(-\frac{\big(x-(x_{p}+\frac{x_{i}}{2})\big)^{2}+\big(y-(y_{p}+\frac{y_{i}}{2})\big)^{2}}{2B_{h}^{2}}\Bigg)\,$, (7)
In the formulas above, $\mathbf{r}_{p}=(x_{p},y_{p})$ is the coordinate of the center of the nucleon in the transverse plane, $\kappa=0.5$ is the gluon fraction and $B_{h}=0.17\text{ fm}$ is the hot-spot size. All these parameters were fixed in previous works nosso ; deb20 ; schenke ; disserta ; glazek-p .
@@
$=$ $\frac{(1-\kappa)^{2}}{36\pi B_{h}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i})-(x_{a}+x_{j})\big)^{2}+\big((y_{p}+y_{i})-(y_{a}+y_{j})\big)^{2}}{4B_{h}^{2}}\bigg)\,$,
@@
$=$ $\frac{\kappa^{2}}{36\pi B_{h}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\Bigg(-\frac{\big(b+(x_{p}+x_{i}/2)-(x_{a}+x_{j}/2)\big)^{2}+\big((y_{p}+y_{i}/2)-(y_{a}+y_{j}/2)\big)^{2}}{4B_{h}^{2}}\Bigg)\,$,
@@
$=$ $\frac{\kappa(1-\kappa)}{36\pi B_{h}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i}/2)-(x_{a}+x_{j})\big)^{2}+\big((y_{p}+y_{i}/2)-(y_{a}+y_{j})\big)^{2}}{8B_{h}^{2}}\bigg)\,$,
@@
$=$ $\frac{\kappa(1-\kappa)}{36\pi B_{h}^{2}}\sum_{i=1}^{3}\sum_{j=1}^{3}\exp\bigg(-\frac{\big(b+(x_{p}+x_{i})-(x_{a}+x_{j}/2)\big)^{2}+\big((y_{p}+y_{i})-(y_{a}+y_{j}/2)\big)^{2}}{8B_{h}^{2}}\bigg)\,.$
@@ II.1.4 Baryon Junction BJ2
The baryon junction BJ2 is a numerical Ansatz for the Y-like structure, as employed in schenke . In this approach we draw the position of the three quarks in the 3D space according to the distribution:
@@ II.2 Initial condition for the nucleus
where $\rho_{0}$ is the nucleon density in the center of the nucleus, $R$ is the radius of the nucleus, $w$ represents deviations from the sphere format, and $a$ is the skin depth. Some usual values for these parameters can be found in Pbparam . We perform our simulation by sorting 208 nucleon center positions, $\mathbf{r}_{n}$, and then inserting the chosen nucleon distribution around it. In the end, the Pb thickness is:
@@ II.3 $k_{T}$-factorization
$\phi_{KLN}(\mathbf{k},x)=\frac{2C_{F}}{3\pi^{2}}\frac{(1-x)^{4}}{\alpha_{s}}\left\{\begin{matrix}1,&\text{, if }k\leq Q_{s}\\ \big(Q_{s}/k\big)^{2}&\text{, if }k>Q_{s}\end{matrix}\right$. (19)
@@
$Q_{s}^{2}(x,\mathbf{r_{\perp}})=T(\mathbf{r_{\perp}})\,\,\frac{2\text{ GeV}^{2}}{1.53\text{ fm}^{-2}}\,\,\bigg(\frac{0.01}{x}\bigg)^{\bar{\lambda}}\,$, (20)
@@
The MC-KLN convolutes the initial conditions of p and p (Pb) with the $k_{T}$-factorization formula of Eq.(17) by calculating $dN/dy\,d^{2}\mathbf{r_{\perp}}$ of gluons at a given rapidity in the overlap region. One example of each initial state geometry for pPb collisions is shown in Fig.2, for the same seed. On the left column we present the proton thickness, on the middle column we show the thickness of the lead nucleus and on the right column we show the convolution of the initial conditions and Eq.(17) at midrapidity.
In Fig.2 we can observe that, remarkably, the produced gluon distribution on the right column inherits the proton shape, which makes pPb collisions good systems to probe the proton initial geometry. This happens because the proton size, which is about $1\text{ fm}$, is much smaller than the lead radius ($6.62\text{ fm}$). Thus, the overlap region will invariably present the proton shape. On the other hand, in pp collisions, since the projectile and target are comparable in size, only fragments of the proton structure will overlap. More specifically, as represented in Fig.1, for hard-sphere and Gaussian initial conditions the convolution will only inherit the proton structure for $b=0$ events, and for the baryon junctions, only for specific angle combinations. Since these conditions are rare, in general this system become less sensitive to the initial proton geometry than pPb.
@@ II.4 Intrinsic fluctuations
So far we have considered only geometrical fluctuations, i.e., those arising from changing positions of the initial quarks and gluons in the transverse plane. In order to describe high-multiplicity events, we must include dynamical fluctuations in the saturation scale larry16 of a given nucleon in the collision. These fluctuations are characterized by deviations of the saturation scale from its mean value and they follow the distribution:
@@ III Results
Integrating Eq.(17) in $p_{T}$, in $y$ (or $\eta$) and in $r_{\perp}$ in an event-by-event basis, we construct the probability $P(N_{ch})$ of producing $N_{ch}$ charged particles. The results for pp and pPb collisions are shown in Fig.3 and Fig.4, respectively. In the figures $K$ is the multiplicative constant of Eq.(17). For fixed values of $K$ and $\sigma$ we integrate the $k_{T}$ factorization formula and obtain values of $N_{ch}$. Accepting all the values of $K$ and $\sigma$ which yield values of $N_{ch}$ which are consistent with the experimental values (within the errors) and which are quoted in the figures, leads to sets of values which define the areas shown in the figures.
The $k_{T}$ factorization formalism is appropriate to describe inelastic non-diffractive events which yield any number of produced particles above a certain value, which we choose to be $N_{ch}=3$. In terms of the KNO variable, this means something like $N_{ch}/\langle N_{ch}\rangle\geq 0.5$. Another limitation of our formalism is that we did not include the contribution from the valence quarks. They are supposed to be important only in the large rapidity regions, where one can use the hybrid formalism duja . At the low rapidities considered here their contribution should be less important but still needed for a better agreement with data.
@@
Results for $P(N_{ch})$ are often presented in terms of the so called Koba-Nielsen-Olsen (KNO) scaling Koba:1972ng variables. KNO scaling means that the multiplicity distribution, $P(N_{ch})$, becomes energy independent at asymptotically high energies when we multiply it by the mean value $\langle N_{ch}\rangle$ and plot it as a function of the scaling variable $z=N_{ch}/\langle N_{ch}\rangle$. This feature has been experimentaly observed at the LHC for narrow pseudo-rapidity ($|\eta|<0.5$) intervals CMS:2010qvf ; at the same time, the violation of KNO scaling has been reported for larger pseudo-rapidity intervals.
In order to reconcile the CGC formalism with KNO scaling, it has been shown that the theory must be nearly Gaussian (corresponding to the strong color field limit) and running coupling effects must be included as well Dumitru:2012tw . Such result was only verified for pp collisions up to $\sim 3.5$ times the average multiplicity. Small violations of the KNO scaling have been reported for pPb collisions in the region where $N_{ch}\lesssim 3$, where additional sources of fluctuations (geometric) are present Dumitru:2012yr . Here, we extend the investigation to larger multiplicities. It is worth noting that KNO scaling might need to be modified for very large multiplicity events Moriggi:2025qfs . In spite of all these efforts, a clear connection between KNO scaling and QCD is not yet established. For complemetary references on this subject we refer the reader to henrique ; germano1 ; germano2 .
The first measurements of multiplicity distribution in pPb collisions alice24 were taken at two different energies. It is possible to make a first check of the validity of the KNO, which, as can be seen in Fig.5 a, seems to be valid. In that figure we see that both experimental data and the MC-KLN model with BJ1 initial geometry exhibit KNO scaling.
Using the BJ1 initial condition we estimate the multiplicity distributions for different pseudorapidity bins. In Fig.5 b we show our predictions for the pseudorapidity ranges $|\eta|<0.5$ and $|\eta|<1.0$. We also provide the mean value estimate and observe that KNO scaling is still valid for these pseudorapidity intervals.
Looking at Fig.3 we observe that data from pp collisions are better described by Gaussian initial conditions, except for the region $z=N_{ch}/\langle N_{ch}\rangle\geq 4$, where BJ1 and BJ2 initial conditions show a better agreement with the high multiplicity data. From Fig.4 we see that pPb data are better described by BJ1. A remarkable feature is that the results from BJ1 and BJ2 are close to each other (within uncertainties) in pp collisions but are quite different in pPb collisions. In the low multiplicity region almost all curves slightly underpredict the data. In Figs.4b and 4d we see the crucial role played by intrinsic fluctuations at large values of $N_{ch}$.
To summarize, we would like to make the following remarks: first, from Figs.1 and  2 we conclude that pPb collisions are better places to study the initial proton geometry than pp collisions; second, our model can be used to calculate the measured multiplicity distributions and the results are sensitive to the initial state inputs. Further constraining the model parameters and reducing the theoretical error, the results would be different enough to be discriminated by the data, which would “prefer” one initial condition. For the moment, in pPb and ultracentral pp collisions ($z\geq 4$) the data favor the baryon junction models. Moreover, in pPb collisions it is possible to discriminate between differents baryon junction models and BJ1 is clearly favored. This is a consequence of the high sensitivity of pPb to the proton initial condition.
@@
Figure 5: a) KNO plot for $\sqrt{s}=5.02$ and $8.16\text{ TeV}$. The experimental data are from alice24 ; b) Predictions for pPb multiplicity distributions at $\sqrt{s}=5.02\text{ TeV}$ for two rapidity intervals $|\eta|<0.5$ and $|\eta|<1.0$ presented in the KNO form.
IV Conclusion
This work is part of a broader project of looking for experimental manifestations of the baryon junction configuration of the proton. Here we investigated proton-proton and proton-nucleus collisions at the LHC. More specifically, we calculated the multiplicity distributions measured in pp and pPb collisions. We used different initial state geometries as input in the MC-KLN Monte Carlo event generator which implements the $k_{T}$-factorization formalism of the CGC with KLN unintegrated gluon distributions. We found that in pp collisions the Gaussian spatial configuration is favored, except in the region $z\geq 4$, where BJ1 and BJ2 give a better description of the data. In pPb collisions the data are better explained by BJ1 initial conditions, provided that fluctuations in the saturation scale (also called “intrinsic fluctuations”) are included.